Aug 08, 1999 OWL shift summary

Richard Kessler (rkessler@fnal.gov)
Sun, 8 Aug 1999 07:08:08 -0500 (CDT)

Shift Summary for OWL shift, Aug 08, 1999

On Shift: RSK and MB

Hours of beam delivered at all intensities: 7
Hours of beam delivered at nominal intensity: 7
Hours of E832 data (special runs not included): 1.5
Hours of special runs (laser scans, muon runs...): 10 min BEAM_MU_4E12

Accounting of lost time:
o identifying and swaping bad divider for CsI 1137
o idem later in the shift.
o swap divider for csi 1542.

========================================================================
Runs ended this shift:
Run 14138 - 9960 VAC00 - BEAM_832 (but last 1/4 is no good)
Run 14140 - 800k events - BEAM_MUON
Run 14142 6941 VAC00 - BEAM_832 (but last 1/4 is no good)

========================================================================
Accesses during shift:
two: to swap csi dividers

========================================================================
Overview of shift: CsI 1137 dropped out around spill 90 of
run 14138, shortly after the OWL shift began.

After a brilliant divider swap, we took good data again for
80 minutes until the same channel [1137] dropped out again from
the laser illumination. We switched HV at the MF with its neighbor
to verify that the problem is at the DPMT end and not the MF HV.
However, before doing this we accidentally swapped two different
channels [1543<-->1542], and after switching them back chan 1542
only holds 1000 V instead of its demand value of 1147 V. Chan 1543
is OK at 1592V, which means that chan 1542 was briefly exposed to
an extra 450V ... which seems to have affected 1542 [right between
the beam holes].

At the end of this shift 1542 divider was swapped and holds HV.
Chan 1137 divider is currently being swapped after Ron checked
the the correct HV arrives at the rib.

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
14138 ~90 00:05 CsI 1137 dropped out of a laser illumination.
Julie notified.

14139 01:00 Short BEAM_832 run to verify that channel 1137 anode
is dead as well as a dynode. Conclude divider is bad.
Julie is on the way.

14140 01:30 BEAM_MU_4E12 while waiting for Julie to come in.
Magnet polarity was swapped to positive.

01:50 Rick & Julie went down to swap bad divider channel
1137.

14141 02:40 Returned from access and short COSMIC run shows that
everything is OK.

14142 01 02:55 Started taking BEAM_832 data again.
Finally have time to read about the
DZero Fiber Tracker Upgrade.

4:10 Noticed high-side tail in the event size distribution
of csi stream 2. See XXI.03.

59 04:15 Pipeline errors. Fixed by usual pipe.tcl commands.
60 CsI 1137 is dead again !!! Paged Julie and BobT.

4:45 End run to debug csi 1137.

========================================================================
Follow up Notes on Previous Shift

Did muon run + polarity swap.

========================================================================
Items to Follow up on Next Shift

- need to remove a few monitor channels from csimon832.nml
to prevent eventual hbook memory overflow.

- ktevana errors seem to always occur at start of run.
We did a kfrend before hitting the START button, but
still got ktevana errors (from crate 53).